OpenStudy (anonymous):

1.0 g of sodium hydroxide, NaOH, was added to 99.0 g of water. The temperature of the solution increased from 18.0 ºC to 20.5 ºC. The specific heat capacity of the solution is 4.18 J g–1 K–1. Which expression gives the heat evolved in kJ mol–1?

sam (.sam.):

Using \[Q=mc \Delta T\] Q=(99.0)(4.18)(2.5) Q=1034.55J 1.0 g of sodium hydroxide=0.0435mol Na Then use \[\Delta H={Q \over n}\] And get your answer \[\Delta H =\frac{1034.55}{0.0435}\]

sam (.sam.):

oh I didn't see that theres a mistake \[1gNaOH=\frac{1molNaOH}{40gNaOH} \times \frac{1molNa}{1molNaOH} =0.25molNa\] Then use \[\Delta H={Q \over n}\] And get your answer \[\Delta H =\frac{1034.55}{0.25}\] Thanks for catching @Callisto :)